1. Solve for the variable
ln ( e^x / e^2x+1 )= -x-1

Answer by ankor@dixie-net.com(22740)

ln ( e^x / e^2x+1 )= -x-1
:
Which is:
ln(e^x) - ln(e^(2x+1)) = -x - 1
:
x*ln(e) - (2x+1)*ln(e) = -x - 1; log equivalent of exponents
:
x*1 - (2x+1)*1 = -x - 1; remember that ln(e) = 1
:
x - 2x - 1 = -x - 1
:
-x - 1 = -x - 1
:
This indicates that x can be any value